OpenStudy (anonymous):

What is the solution of the system of equations? y = 3x + 7 y = x – 9

OpenStudy (anonymous):

(-4,-13)

OpenStudy (mrnood):

@LilStacie Giving the answer like that is a) Not permitted in the code of conduct b) not helpful to the poster - she has no way of knowing that oyu are correct and learns NOTHING from copying your answer @siahniya16 you jave 2 equations both telling you y=? the solution is when both those equations are true so you can write y = 3x+7 = x-9 or 3x+7 = x-9 can you work out what x is from that?

OpenStudy (mrnood):

@lilstacie you answer is not correct either!

OpenStudy (mrnood):

you have 2 equations both telling you y=? the solution is when both those equations are true so you can write: y = 3x+7 = x-9 or 3x+7 = x-9 can you work out what x is from that?

OpenStudy (texaschic101):

just do : 3x + 7 = x - 9 and solve for x. Once x is found, sub that back into either of the original equations to find y